Find the slope of the line:
a. through (5, -2) and (-1,-9)
b. with equation 4x -5y = 10

Answer by Tutorteddy.Com(12)

a. slope = [(-9 -(-2)]/(-1 - 5) = (-7)/(-6) = 7/6
b. slope = -4/(-5) = 4/5.